OpenStudy (anonymous):

Which is the graph of the equation y – 2 = 0.5(x + 3)?

OpenStudy (anonymous):

y - 2 = (-1/2)x - 1.5 y = (-1/2)x + 0.5 slope = -1/2 y-intercept = 0.5 (0, -0.5) x-intercept: 0 = (-1/2)x + 0.5 0.5x = 0.5 x = 0.5/0.5 = 1 x -intercept = (1,0)